• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 07:58:43 on February 11, 2007

    Entry number 192294

    keyword=2/11 Owl Shift Log

    00:00 Shift starts on tantalum target with Alexander working to understand higher than expected dead times. Hrs-L angle 41.5 deg. Kinematics setup E.

    00:12 Request 10uA from MCC.

    00:28 Requested Controlled Access to look into the high dead-time issues.

    00:53 Hall C requests the half-wave plate be removed. Next run 6181.

    01:35 Request Beam Permit

    01:45 MCC called to ask what current we would like to be restored to. Requested 15 uA.

    01:50 Called MCC back to get 15uA rather than 10uA. Our greed, caused and FSD...

    03:56 Finished with Ta target & requested move to Carbon.

    03:58 Completed target move and requested return of beam at 5uA. Begin production running on carbon target.

    04:47 Completed carbon target production running & requested change of target to boron carbide.

    04:49 Completed target move & requested 5uA beam from MCC. MCC is doing their QE.

    04:56 Began taking proudction data on B4C target.

    05:26 Completed B4C target & requested move to Li target.

    05:31 Completed move to Li target & requested 5uA from MCC.

    06:10 Completed Li production running, requested target change to Ta, and moved spectrometer (HRS-L) to 46.0 degrees (readback 46.003 degrees).

    06:20 Request 15uA beam from MCC to commence production running on Ta with new kinematics (Kinematics F).

    07:42 Please note, at least with Thallium, we have been doing 1hr runs in lieu of 1M event runs per instructions on white board/bulletin board in counting house.

    07:58 Shift concluded on Ta target.